| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  114  
 | 
Sun May  9 21:29:26 2021 | 
Helmut, Wolfram, Yuri | Accelerator | target TE3UF shift | We see a very strong line at the left end 
of spectrum, stronger than all other lines.
 
When we shift the target TE3UF_H from -35mm | 
   | 
| 
  113  
 | 
Sun May  9 20:39:17 2021 | 
Shahab | Runs | Subtracting the background | still in the search of 70-Se | 
    | 
| 
  112  
 | 
Sun May  9 20:34:47 2021 | 
Rui-Jiu Chen | Analysis | Comparison of  frequency spectrum between 72Ge and other isotope | This some preliminary analysis results by 
using  NTCAP data.   The IQ rate of NTCAP 
is 20 MSamples/seconds. I analyzed data files | 
 6x   | 
| 
  111  
 | 
Sun May  9 19:03:02 2021 | 
Rodostina, Oliver, Wolfram, Helmut, Yuri, Shahab | Runs | quadrupoles were down, fixed. |   | 
   | 
| 
  110  
 | 
Sun May  9 18:50:23 2021 | 
Helmut, Wolfram, Yuri | Accelerator | changed to 70Se | 18:45  changed SIS, ESR setting for 70Se without 
predrilling blind change.
 
E-SIS =  446.55 MeV/u, is centerd as before | 
   | 
| 
  109  
 | 
Sun May  9 18:41:33 2021 | 
Rodostina, Oliver, Wolfram, Helmut, Yuri, Shahab | Runs | stopped data taking for energy change | last recorded filename:
 
 
E143-245MHz-2021.05.09.17.38.44.410.tiq
 | 
   | 
| 
  108  
 | 
Sun May  9 17:39:21 2021 | 
Guy | Analysis | Comparison between 72Ge: 7th 19h vs 8th 21h | Following Xiangcheng Chen's analysis, I analysed 
11 ms time slices separated by 25 ms to see 
the evolution of the potential isomer across | 
     | 
| 
  107  
 | 
Sun May  9 16:19:06 2021 | 
Rodostina, Oliver, Wolfram, Shahab | Runs | Setting to 52Mn ???? | first filename in this recording:
 
 
E143-245MHz-2021.05.09.15.16.49.816.tiq | 
   | 
| 
  106  
 | 
Sun May  9 14:19:44 2021 | 
Yuri | Accelerator | Spectrum | We accept larger m/q values than we like to. 
m/q(low) cut ~ 2.08. Decrease ESR energy 
from 369.2 -> 368.4 MeV/u
 | 
   | 
| 
  105  
 | 
Sun May  9 13:53:07 2021 | 
Yuri, Wolfram, Andrew, Liliana | Calibration | Change of energy: Back as before | After looking at the spectrum it was concluded 
that the same isotopes, and no more, could 
be identified. 
 | 
    | 
| 
  104  
 | 
Sun May  9 13:15:38 2021 | 
Liliana Andrew | Runs | Stopped data taking during modifications for the beam energy | Beam energy to be increase to move Mn into 
the ring. 
 
First step we set the energy to 445 MeV. | 
    | 
| 
  103  
 | 
Sun May  9 12:53:01 2021 | 
Desislava Kalaydjieva, Taka Yamaguchi | Analysis | Analysis, Morning shift, 9 May | For the whole morning shift there were no 
significant changes in the spectra.
 
 | 
 9x   | 
| 
  102  
 | 
Sun May  9 08:00:03 2021 | 
Liliana, Andrew | General | Status  | So far, we keep the setting as it is and wait 
for experts to come in. 
 
Few spills are missing from time to time. | 
    | 
| 
  101  
 | 
Sun May  9 07:16:47 2021 | 
Iris, Chris G., Jacobus, Michele | Analysis | Runs 3-4am Not_Mn52 | Summed up runs on the 245 MHz detector between 
3-4am (computer time).
 
-s 0.1 -t 8.0 | 
     | 
| 
  100  
 | 
Sun May  9 06:53:20 2021 | 
Iris, Chris G., Jacobus, Michele | General | Summary night shift | The tune last night was unfortunately not 
to 52Mn. The two "isomers" that were identified 
are instead 45Sc and 75Br. But separation | 
   | 
| 
  99  
 | 
Sun May  9 06:34:37 2021 | 
Iris, Chris G., Jacobus, Michele | Analysis | Runs 2-3am Not_Mn52 | Summed up runs on the 245 MHz detector between 
2-3am (computer time).
 
One has -s 0.1 -t 8.0
 | 
       | 
| 
  98  
 | 
Sun May  9 06:15:09 2021 | 
Iris, Chris G., Jacobus, Michele | Runs | Triplet: zoom in | Zoom into the triplet: the satellite peaks 
on the right are clearly visible.
 
This is just one file but one *might* see | 
    | 
| 
  96  
 | 
Sun May  9 05:56:29 2021 | 
Iris, Chris G., Jacobus, Michele | Analysis | Not_52Mn spectra added up | Attached is a summed spectra of all files 
taken between 1:17am and 1:59am (on the 245 
MHz detector).
 | 
     | 
| 
  95  
 | 
Sun May  9 05:54:11 2021 | 
Ragan | Analysis | offset btw png and tiq files from server? | This is due to the 50% trigger setting in 
the spectrum analyzers )) | 
   | 
| 
  94  
 | 
Sun May  9 05:38:29 2021 | 
Iris, Chris G. | Analysis | offset btw png and tiq files from server? | This might have been answered before but I 
looked at one png file and one tiq file from 
the same run that was stored on the server. | 
    |